# Online Learning of Whittle Indices for Restless Bandits with Non-Stationary Transition Kernels

### v1

@@ 6 Numerical Study
10 4 426.46±$\pm 23$ 958.96±$\pm 63$ 4579.8±$\pm 39$ 5172.2±$\pm 37$
20 4 656.53±$\pm 41$ 2637.1±$\pm 155$ 13609±$\pm 178$ 12584±$\pm 137$
Table 1: $\mathrm{Reg}(T)$ for $T=50$ episodes for different values of $N$ and $M$ with $H=100$ and $\epsilon_{n}=0.05$
@@
In this section, we demonstrate the performance of our proposed policy by evaluating it under two applications modeled as restless multi-armed bandit problems. In each application, we consider that there are $N$ arms and a policy can activate $M$ of them in each time slot $h\in[H]$ of every episode $t\in[T]$. We evaluate our policy against the UCWhittle policy [23], the WIQL policy [27], and a uniformly randomized policy [3]. The resultsare averaged over $50$ independent runs.
One Dimensional Bandits: We consider a modified version of the one dimensional RMAB problem studied in [38, 33]. Each arm $n$ is a Markov process with $K$ states, numbered as $10,1,\ldots,K-$ 1. For our simulations, we set $K=10$. The reward of an arm increases linearly with the current state, i.e. $r(s,a)=s$. If the arm is activated, then it can evolve from state $s$ to $\min\{s+1,K-1\}$ with probability $q_{n}(t)$ or remain in the same state $s$ with probability $1-q_{n}(t)$. If the arm is not activated, then it evolves from state $s$ to $\max\{s-1,0\}$ with probability $p_{n}(t)$ or remain in the same state $s$ with probability $1-p_{n}(t)$. One-dimensional MDPs of this form are often used in health monitoring and machine monitoring applications [39, 40]. In our simulation, for half of the $N$ arms, we consider the following kernel dynamics which are unknown and must be learned from data - (i) $p_{n}(t)$ changes to $\min\{p_{n}(t-1)+\epsilon_{n},1\}$ with probability $0.7$, or it changes to $\max\{p_{n}(t-1)-\epsilon_{n},0\}$ with probability 0.3 and (ii) $q_{n}(t)$ is unknown but stationary. For the other half, probabilities $p_{n}(t)$ and $q_{n}(t)$ are known to the decision maker and do not change over episode $t$. The detailed experimental setup is provided in the Appendix D.1.
Wireless Scheduling: We consider a wireless scheduling problem, where $M$ out of $N$ sources can send their observation to a receiver side over an unreliable channel at every time slot $h\in[H]$ of episode $t\in[T]$. Due to channel unreliability, the observation may not be delivered. The goal of the receiver is to estimate the current signal values of all $N$ sources based the information delivered from the sources. The reward for accurate timely estimation can be modeled as the mutual information between the estimated signal and the actual signal. [41] showed that the mutual information can be determined by using a decreasing function $-(\mathrm{log}_{2}(1-\sigma_{n}^{2s_{n,h,t}}))/2$ of Age of Information (AoI) for zero-mean i.i.d. Gaussian random variables with variance $\sigma^{2}_{n}$, where AoI $s_{n,h,t}$ of source $n$ is the time difference between current time $h$ and the generation time of the most recently delivered signal. The AoI value of a source $n$ increases by $1$ if the source $n$ is not scheduled. If the source $n$ is scheduled, the AoI value drops to $1$ with probability $q_{n}(t)$ (successful delivery) or increases by $1$ with probability $1-q_{n}(t)$ (unsucessful delivery). The parameter $q_{n}(t)$ measures the reliability of channel $n$ at time $t$. In our experiment, we assume that $q_{n}(t)$ is unknown and non-stationary for half of the sources, whereas it is is unknown but stationary for the remaining half. The detailed experimental setup is provided in the Appendix D.2.
Figure 1: One Dimensional Bandit: Cumulative regret versus the number of episodes with different values of time horizon ($H=20,50,100$) with $\epsilon_{n}=0.05$.
Figure 2: Wireless Scheduling: Cumulative regret versus the number of episodes with different values of time horizon ($H=20,50,200$) with $\epsilon_{n}=0.05$.
Simulation results: Simulation results are shown in Table 1, Figure 1 and Figure 2. It can be observed that our algorithm achieves the best regret in all cases. UCWhittle [23] is designed for static settings, hence it uses all historical data to learn transitions, which often yields inaccurate transition kernels for the current episode. Moreover, UCWhittle naively attempts to learn the entire transition matrix. The stationary random policy [3] picks $M$ out of $N$ arms uniformly at random, clearly leading to worse performance. WIQL [27], a Q-learning approach, requires many data samples to converge. Also, similar to UCWhittle, WIQL uses all historical data. Due to WIQL’s slow convergence, WIQL shows similar performance compared to random policy, see [23]. Because of the non-stationary setting, WIQL performs poorly even compared to random policy, suggesting that traditional reinforcement learning methods cannot be naively applied to solve dynamic RMAB problems.
@@ 7 Conclusions and Limitations
This paper introduced an online/reinforcement learning algorithm for estimating the Whittle index for restless bandit problems with unknown and non-stationary transition kernels using sliding window and upper confidence bound approaches. To our knowledge, this is the first work to provide an upper bound of the dynamic regret of an online Whittle index-based algorithm for RMABs with unknown and non-stationary transition kernels. Our proposed algorithm is evaluated on two different restless bandit problems against three baselines and provides significant performance gains. We also provide novel regret analysis and tighter regret bounds for RMABs in both stationary and non-stationary settings. An interesting direction of future work involves proving lower bounds for regret, i.e., showing that no online algorithm can achieve regret better than $O(\sqrt{T\mathrm{log}(T)})$. Moreover, our algorithm can only be applied to a finite state space and works best for small $\epsilon_{n}$ (slowly varying kernels). Promising future directions include extending this work to infinite or continuous state spaces, and designing algorithms that achieve sub-linear dynamic regret even for large $\epsilon_{n}$ (rapidly varying kernels).
Appendix A Proof of Lemma 1
We can get Lemma 1 by combining Lemma 2 and Lemma 3, which are provided below:
Lemma 2.
Given $\eta_{1}>0$ and $t\geq 1$, for $(s,a)\in\mathcal{Z}^{(n)}_{1}$, we have
$\mathrm{Pr}\left(\|\hat{P}_{n,t,t-1}(\cdot|s,a)-P_{n,t}(\cdot|s,a)\|_{1}\geq\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{1}^{(n)}|NT/\eta_{1})}{C^{(n)}_{t,t-1}(s,a)}}\right)\leq\frac{\eta_{1}}{N|\mathcal{Z}^{(n)}_{1}|T}.$ (20)
Proof.
@@ Appendix A Proof of Lemma 1
$\mathrm{Pr}(|\hat{p}-p|_{1}\geq\beta)\leq(2^{m}-2)\exp^{(-\frac{n\beta^{2}}{2})},$ (21)
where $n$ is the number of samples.
By applying (21) with
$\beta=\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{1}^{(n)}|NT/\eta_{1})}{C^{(n)}_{t,t-1}(s,a)}}$
@@
$n=C^{(n)}_{t,t-1}(s,a),$
we have
$\mathrm{Pr}\left(\|\hat{P}_{n,t,t-1}(\cdot|s,a)-P_{n,t}(\cdot|s,a)\|_{1}\geq\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{1}^{(n)}|NT/\eta_{1})}{C^{(n)}_{t,t-1}(s,a)}}\right)$
$\leq 2^{|\mathcal{S}|}\exp(-\mathrm{log}(2^{|\mathcal{S}|}|\mathcal{Z}_{1}^{(n)}|NT/\eta_{1}))=\frac{\eta_{1}}{N|\mathcal{Z}_{1}^{(n)}|T}.$ (22)
Lemma 3.
Given $\eta_{2}>0$ and $w\geq 1$, for $(s,a)\in\mathcal{Z}_{2}$, we have
$\mathrm{Pr}\left(\|\hat{P}_{n,t,w}(\cdot|s,a)-P_{n,t}(\cdot|s,a)\|_{1}\geq\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{2}^{(n)}|NT/\eta_{2})}{C^{(n)}_{t,w}(s,a)}}+w\epsilon\right)\leq\frac{\eta_{2}}{N|\mathcal{Z}^{(n)}_{2}|T}.$ (23)
Proof.
By using (21) with
$\beta=\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}^{(n)}_{2}|NT/\eta_{2})}{C^{(n)}_{t,w}(s,a)}}$
and
$n=C^{(n)}_{t,w}(s,a),$
we have
$\mathrm{Pr}\left(\|\hat{P}_{n,t,w}(\cdot|s,a)-\mathbb{E}[\mathbf{1}(\cdot,s,a,n,t,w)]\|_{1}\geq\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}^{(n)}_{2}|NT/\eta_{2})}{C^{(n)}_{t,w}(s,a)}}\right)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq\frac{\eta_{2}}{N|\mathcal{Z}^{(n)}_{2}|T}.$ (24)
@@
$=$ $\|P_{n,t}(\cdot|s,a)-\max_{t^{\prime}\in\{t-w+1,t-(w-1)+1,\ldots,t\}}P_{n,t^{\prime}}(\cdot|s,a)]\|_{1}\leq w\epsilon.$ (25)
By combining (A) and (A), we obtain Lemma 3. ∎
Now, by combining Lemma 2 and Lemma 3, we have
$\mathrm{Pr}(P_{n,t}\in B_{n,t},\forall n,\forall t)$ $\geq 1-\sum_{t=1}^{T}\sum_{n=1}^{N}\sum_{(s,a)\in\mathcal{Z}^{(n)}_{1}}\frac{\eta_{1}}{N|\mathcal{Z}_{1}^{(n)}|T}-\sum_{t=1}^{T}\sum_{n=1}^{N}\sum_{(s,a)\in\mathcal{Z}^{(n)}_{2}}\frac{\eta_{2}}{N|\mathcal{Z}_{2}^{(n)}|T}$
$=1-\eta_{1}-\eta_{2}.$ (26)
@@
Appendix B Proof of Theorem 1
When the confidence bound holds, we have
$\mathrm{Reg}^{(t)}$ $=\sum_{n=1}^{N}U(\pi^{*}_{n,t},P_{n,t},\lambda)-U(\pi_{n,t},P_{n,t},\lambda){\leq}\sum_{n=1}^{N}U(\pi_{n,t},\tilde{P}_{n,t},\lambda)-U(\pi_{n,t},P_{n,t},\lambda)$
$\leq\sum_{n=1}^{N}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}}\gamma^{h-1}(\tilde{P}_{n,t}(s^{\prime}|s_{n,t,h},a_{n,t,h})-P_{n,t}(s^{\prime}|s_{n,t,h},a_{n,t,h}))V_{n}(s^{\prime};\pi_{n,t},P_{n,t})\bigg{]},$
@@ Appendix B Proof of Theorem 1
$=\sum_{n=1}^{N}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{Z}^{n}_{1}}\alpha_{t}^{(n)}(s,a)d_{t,1}^{(n)}(s,a)+\sum_{(s,a)\in\mathcal{Z}^{(n)}_{2}}\alpha_{t}^{(n)}(s,a)d_{t,2}^{(n)}(s,a)\bigg{]}V_{max},$ (27)
where the simplified notation $V_{n}$ is used instead of $V_{n,\lambda,t}$, $V_{max}=\max_{n\in[N],s\in\mathcal{S}}V_{n}(s^{\prime};\pi_{n,t},P_{n,t})$ and $\alpha_{t}^{(n)}(s,a)$ is a random variable denoting the number of visits of $(s,a)\in\mathcal{S}\times\mathcal{A}$ at episode $t\in[T]$.
By substituting the value of $d_{t,1}^{(n)}(s,a)$, we have
$\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{Z}_{1}^{(n)}}\alpha_{t}^{(n)}(s,a)d_{t,1}^{(n)}(s,a)\bigg{]}$
$\leq\sqrt{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}_{1}^{(n)}|NT/\eta_{1})}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{Z}_{1}}\frac{\alpha_{t}^{(n)}(s,a)}{\sqrt{C^{n}_{t,t-1}(s,a)}}\bigg{]}.$ (28)
Moreover, by applying [23, Lemma E.3], we get
$\sum_{t=1}^{T}\sum_{(s,a)\in\mathcal{Z}_{1}}\frac{\alpha_{t}^{(n)}(s,a)}{\sqrt{C^{(n)}_{t,t-1}(s,a)}}\leq(\sqrt{H+1}+1)\sqrt{|\mathcal{Z}_{1}^{(n)}|TH}$ (29)
Hence,
$\sum_{t=1}^{T}\sum_{n=1}^{N}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{Z}_{1}^{(n)}}\alpha_{t}^{(n)}(s,a)d_{t,1}^{(n)}(s,a)\bigg{]}V_{\max}\leq O\left(\bigg{(}\sum_{n=1}^{N}|\mathcal{Z}^{(n)}_{1}|\bigg{)}H\sqrt{T\mathrm{log}T}\right).$ (30)
On the other hand, by substituting the value of $d_{t,2}^{(n)}(s,a)$, we have
$\sum_{t=1}^{T}\sum_{n=1}^{N}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{Z}_{2}^{(n)}}\alpha_{t}^{(n)}(s,a)d_{t,2}^{(n)}(s,a)\bigg{]}$
$=\sum_{t=1}^{T}\bigg{(}\sum_{n=1}^{N}\sqrt{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}^{(n)}_{2}|NT/\eta_{1})}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg{[}\sum_{(s,a)\in\mathcal{Z}_{2}}\frac{\alpha_{t}^{(n)}(s,a)}{\sqrt{C^{(n)}_{t,w_{n}}(s,a)}}\bigg{]}+w_{n}\epsilon_{n}|\mathcal{Z}^{(n)}_{2}|\bigg{)}$
$=\sum_{t=1}^{T}\bigg{(}\sum_{n=1}^{N}\sqrt{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}^{(n)}_{2}|NT/\eta_{1})}\sum_{(s,a)\in\mathcal{Z}^{(n)}_{2}}g_{t,n}(s,a,w_{n})+w_{n}\epsilon_{n}|\mathcal{Z}_{2}|\bigg{)}$
$\leq\sum_{t=1}^{T}\bigg{(}\sum_{n=1}^{N}\sqrt{2|\mathcal{S}|\mathrm{log}(2|\mathcal{Z}^{(n)}_{2}|NT/\eta_{1})}|\mathcal{Z}^{(n)}_{2}|G_{t,n}(w_{n})+w_{n}\epsilon_{n}|\mathcal{Z}^{(n)}_{2}|\bigg{)}$
$=\sum_{t=1}^{T}O\bigg{(}\sum_{n=1}^{N}|\mathcal{Z}^{(n)}_{2}|\bigg{(}G_{t,n}(w_{n})+w_{n}\epsilon_{n}\bigg{)}\bigg{)},$ (31)
where $G_{t,n}(w)=\max_{(s,a)\in\mathcal{Z}^{(n)}_{2}}g_{t,n}(s,a,w)$ and

### v2

@@ 6 Simulation Results
[…]
Table 1: $\mathrm{Reg}(T)$ for $T=50$ episodes for different values of $N$ and $M$.
(a) Wireless Scheduling.
(b) One Dimensional Bandit.
Figure 1: $\mathrm{Reg(T)}$ Vs. number of episodes in Scheduling and 1-D Bandit with $N=20,M=4$.
@@
In this section, we demonstrate the performance of our proposed policy by evaluating it under two applications modeled as RMAB. In each application, we consider that there are $N$ arms and a policy can activate $M$ of them in each time slot $h\in[H]$ of every episode $t\in[T]$. We evaluate our policy against the UCWhittle policy [30], UCWhittle + Window policy, where we incorporate sliding window to UCWhittle and the window size is taken randomly, the WIQL policy [34], and a randomized policy [3]. The results are averaged over $50$ independent runs. Simulation results are shown in Table 1 and Figure 1. It can be observed that our algorithm achieves the best regret in all cases. The performance gain is observed due to two main reasons: (i) we are predicting transition kernels with intelligent update of window size and (ii) our algorithm exploits the knowledge of sparsity. Detailed experimental setup is discussed in Appendix A.5.
@@ 7 Conclusions and Limitations
This paper introduced an online/reinforcement learning algorithm for estimating the Whittle index for restless bandit problems with unknown and non-stationary transition kernels using sliding window and upper confidence bound approaches. To our knowledge, this is the first work to provide an upper bound of the dynamic regret of an online Whittle index-based algorithm for RMABs with unknown and non-stationary transition kernels. Our proposed algorithm is evaluated on two different restless bandit problems against four baselines and provides significant performance gains. We also provide novel regret analysis. An interesting direction of future work involves proving lower bounds for regret. Other future directions include extending this work to infinite or continuous state spaces, and designing algorithms that achieve sub-linear dynamic regret even for large $\tilde{V}_{n}$ (rapidly varying kernels).
Appendix A Appendix
A.1 Proof of Lemma 1
@@
$\mathrm{Pr}(|\hat{p}-p|_{1}\geq\beta)\leq(2^{m}-2)\exp^{(-\frac{k\beta^{2}}{2})},$ (20)
where $k$ is the number of samples.
By using equation 20 with
$\beta=\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{S}||\mathcal{A}|NT/\eta)}{C^{(n)}_{t,w}(s,a)}}$
@@
$k=C^{(n)}_{t,w}(s,a),$
we get
$\mathrm{Pr}\left(\|\hat{P}_{n,t,w}(\cdot|s,a)-\mathbb{E}[\mathbf{1}(\cdot,s,a,n,t,w)]\|_{1}\geq\sqrt{\frac{2|\mathcal{S}|\mathrm{log}(2|\mathcal{S}||\mathcal{A}|NT/\eta)}{C^{(n)}_{t,w}(s,a)}}\right)$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\leq\frac{\eta}{N|\mathcal{S}||\mathcal{A}|T}.$ (21)
@@
$=$ $\|P_{n,t}(\cdot|s,a)-\max_{t^{\prime}\in\{t-w+1,t-(w-1)+1,\ldots,t\}}P_{n,t^{\prime}}(\cdot|s,a)]\|_{1}\leq w_{n}V_{n}/T.$ (22)
Now, by combining equation A.1 and equation A.1, we have
$\mathrm{Pr}(P_{n,t}\in B_{n,t},\forall n,\forall t)$ $\geq 1-\sum_{t=1}^{T}\sum_{n=1}^{N}\sum_{(s,a)\in\mathcal{S}\times A}\frac{\eta}{N|\mathcal{S}||\mathcal{A}|T}$
$=1-\eta.$ (23)
@@
A.2 Proof of Theorem 1
We first decompose $\mathrm{Term1}$. We solve another bandit problem to select $V_{n}$ from a set of possible drift values based on the history by using EXP3 algorithm [41]. In this case, we can decompose the regret associated with $\mathrm{Term1}$ as follows:
$\mathrm{Term1}$ $=\sum_{t=1}^{T}\sum_{n=1}^{N}U(\pi^{*}_{n,t},P_{n,t},\lambda)-U(\pi_{n,t}(\hat{V}_{n}(t)),P_{n,t},\lambda)$
$=\sum_{t=1}^{T}\sum_{n=1}^{N}\bigg(U(\pi^{*}_{n,t},P_{n,t},\lambda)-U(\pi_{n,t}(V_{n}),P_{n,t},\lambda)\bigg)$
$~~+\sum_{t=1}^{T}\sum_{n=1}^{N}\bigg(U(\pi_{n,t}(V_{n}),P_{n,t},\lambda)-U(\pi_{n,t}(\hat{V}_{n}(t)),P_{n,t},\lambda)\bigg)$ (24)
where we denote $V_{n}$ is optimal, $\hat{V}_{n}(t)$ is estimated, and $\pi_{n}(\mathcal{V})$ denotes our policy when $\mathcal{V}$ is used.
The regret bound for the term
$\sum_{t=1}^{T}\sum_{n=1}^{N}\bigg(U(\pi_{n,t}(V_{n}),P_{n,t},\lambda)-U(\pi_{n,t}(\hat{V}_{n}(t)),P_{n,t},\lambda)\bigg)$
represents the loss from needing to learn $V_{n}$ instead of knowing it apriori and can be found by using [41]. In particular, assuming the finite discretization of possible drifts in Section 4.3, we can have
$\sum_{t=1}^{T}\sum_{n=1}^{N}\bigg(U(\pi_{n,t}(V_{n}),P_{n,t},\lambda)-U(\pi_{n,t}(\hat{V}_{n}(t)),P_{n,t},\lambda)\bigg)\leq\sum_{n=1}^{N}O\bigg(\sqrt{J_{n}\mathrm{log}(J_{n})T}\bigg),$ (25)
where $J_{n}$ is the number of elements in the set of possible values of drift for arm $n$.
Next, we show the upper bound of $\sum_{t=1}^{T}\sum_{n=1}^{N}\bigg(U(\pi^{*}_{n,t},P_{n,t},\lambda)-U(\pi_{n,t}(V_{n}),P_{n,t},\lambda)\bigg)$. For the ease of notation, we use $\pi_{n,t}(V_{n})$ as $\pi_{n,t}$. When the confidence bound holds, we have
$\sum_{n=1}^{N}U(\pi^{*}_{n,t},P_{n,t},\lambda)-U(\pi_{n,t},P_{n,t},\lambda){\leq}\sum_{n=1}^{N}U(\pi_{n,t},\tilde{P}_{n,t},\lambda)-U(\pi_{n,t},P_{n,t},\lambda)$
$\overset{a}{=}\sum_{n=1}^{N}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg[\sum_{h=1}^{H}\sum_{s^{\prime}\in\mathcal{S}}\gamma^{h-1}(\tilde{P}_{n,t}(s^{\prime}|s_{n,t,h},a_{n,t,h})-P_{n,t}(s^{\prime}|s_{n,t,h},a_{n,t,h}))V_{n}(s^{\prime};\pi_{n,t},\tilde{P}_{n,t})\bigg],$
@@ A.2 Proof of Theorem 1
$\leq\sum_{n=1}^{N}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg[\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\alpha_{t}^{(n)}(s,a)d_{t}^{(n)}(s,a)\bigg]V_{max}$ (26)
where (a) is obtained by using [30, Theorem 6.4], the simplified notation $V_{n}$ is used instead of $V_{n,\lambda,t}$, $V_{max}=\max_{n\in[N],s\in\mathcal{S}}V_{n}(s^{\prime};\pi_{n,t},\tilde{P}_{n,t})$ and $\alpha_{t}^{(n)}(s,a)$ is a random variable denoting the number of visits of $(s,a)\in\mathcal{S}\times\mathcal{A}$ at episode $t\in[T]$.
By substituting the value of $d_{t}^{(n)}(s,a)$, we have
$\sum_{t=1}^{T}\sum_{n=1}^{N}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg[\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\alpha_{t}^{(n)}(s,a)d_{t}^{(n)}(s,a)\bigg]$
$\leq\sum_{t=1}^{T}\bigg(\sum_{n=1}^{N}\sqrt{2|\mathcal{S}|\mathrm{log}(2|\mathcal{S}||\mathcal{A}|NT/\eta)}\mathbb{E}_{P_{n,t},\pi_{n,t}}\bigg[\sum_{(s,a)\in\mathcal{Z}_{2}}\frac{\alpha_{t}^{(n)}(s,a)}{\sqrt{C^{(n)}_{t,w_{n}}(s,a)}}\bigg]+w_{n}V_{n}/TH\bigg)$
$=\sum_{t=1}^{T}\bigg(\sum_{n=1}^{N}\sqrt{2|\mathcal{S}|\mathrm{log}(2|\mathcal{S}||\mathcal{A}|NT/\eta)}\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}g_{t,n}(s,a,w_{n})+w_{n}V_{n}H/T\bigg)$
$\leq\sum_{t=1}^{T}\bigg(\sum_{n=1}^{N}\sqrt{2|\mathcal{S}|\mathrm{log}(2|\mathcal{S}||\mathcal{A}|NT/\eta)}2|\mathcal{S}|G_{t,n}(w_{n})+w_{n}V_{n}H/T\bigg)$
$=\sum_{t=1}^{T}O\bigg(\sum_{n=1}^{N}2|\mathcal{S}|G_{t,n}(w_{n})+w_{n}V_{n}H/T\bigg),$ (27)
where $G_{t,n}(w)=\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}g_{t,n}(s,a,w)$ and